• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:03:19 on February22,2001

    Entry number 56765

    keyword=cryotarget switch from LH2/LD2 to He4 completed.

    Cryotarget switch from LH2/LD2 to He4:
    Started preparation after 5 am. Completed warmup by 9:30.
    Completed Pump/purge for loops 2 and 3 (H and D), filled both loops
    with gas He4 (by 10:30).
    Completed cooldown of He4 loop by 15:45.
    Loop 1 filled with He4 at 5.8K and 221 psi. Now HP power 50W, Fan at 10% (24 Hz).
    Loop2 filled with He4 at 156.5K at 25.5 psi. HP heater and Fan are off.
    Loop 3 filled with He4 at 147.7K at 24.1 psi. . HP heater and Fan are off.
    Coolant flow before 4K cool-down (spectrometer use) was around 25 g/s.
    Now with minimum power flow is at 33.2 g/s.
    Target switch completed (took about one and half shifts).